# Conformal Quantile Regression for Probabilistic Constitutive Modeling of Anisotropic Soft Materials

### v1

@@ Abstract
Keywords Constitutive Law Modeling $\cdot$ Uncertainty Quantification $\cdot$ Conformal Quantile Regression $\cdot$ Monotonic Neural Network $\cdot$ Thermodynamics Consistency $\cdot$ Scientific Machine Learning
@@ 1 Introduction
In this work, we construct a physics-encoded, anisotropic data-driven constitutive model, building on the seminal framework of [50]. Anisotropy is incorporated through structural tensors that encode the material symmetries associated with the underlying anisotropic directions. Within this formulation, we consider neural network parameterizations for the phenomenological components of the constitutive response that preserve polyconvexity by construction. Rather than parameterizing the unknown strain energy density itself [28, 6], we parameterize its required gradients, which enter directly into the stress response. This choice leads to improved trainability and reduced computational cost.
@@
The enforcement of thermodynamic constraints in a fully probabilistic setting presents an additional challenge. Incorporating physical principles—such as polyconvexity, material stability—into Bayesian neural network frameworks [10] is generally nontrivial and may require specialized parameterizations or carefully designed constrained priors to ensure that both prior and posterior distributions remain physically admissible [11].
@@
1.1 Other Related Work
@@ 2.1 Deterministic Constitutive Backbone
Specifically, the strain energy density is expressed as
$\psi\left(\bm{F};\{\overset{j}{\bm{G}}\}_{j=1}^{N_{G}}\right)=\sum_{i=1}^{2}f_{i}(I_{i})+\sum_{j=1}^{N_{G}}\left[\bar{f}_{4j}(I_{4j})+\hat{f}_{5j}(I_{5j})\right],$ (1)
@@
$I_{4j}=\mathrm{tr}\!\left[\bm{C}\,\overset{j}{\bm{G}}\right],\quad I_{5j}=\mathrm{tr}\!\left[\mathrm{cof}(\bm{C})\,\overset{j}{\bm{G}}\right].$ (2)
@@
The functions $f_{i}$, $\bar{f}_{4j}$, and $\hat{f}_{5j}$ are unknown real-valued scalar functions. If the scalar functions are chosen to be convex with respect to their arguments, the resulting strain energy density is polyconvex in $\bm{F}$ [36, 7]. As a consequence, the associated variational problem admits minimizers under standard boundary conditions, and the resulting constitutive model satisfies rank-one convexity, which is closely related to mechanical stability. Moreover, adherence to this framework guarantees thermodynamic consistency and satisfies the required symmetry properties with respect to the structural tensors (or fiber orientations).
@@
$\psi\!\left(\bm{F};\{\overset{j}{\bm{G}}\}_{j=1}^{N_{G}}\right)=\sum_{k=1}^{M}\tilde{f}_{k}(\tilde{I}_{k}),\qquad M=2+2N_{G}.$ (3)
@@
where $\tilde{g}_{k}(\tilde{I}_{k})\equiv d\tilde{f}_{k}/d\tilde{I}_{k}$ denotes the derivative of the corresponding basis function. Since each $\tilde{f}_{k}$ is convex and non-decreasing, its derivative $\tilde{g}_{k}$ is non-negative and monotone, a property that plays a key role in the subsequent parameterization and learning strategy. For notational simplicity, the incompressibility contribution, which enters as an additional additive term $p\,\mathbf{F}^{-T}$, is not shown explicitly but is included in all computations.
@@ 2.2 Parameterization Strategy
Intuitively, searching over monotone functions is simpler and more stable than searching over convex functions. Once $\tilde{g}_{k}$ has been learned, the corresponding energy functions $\tilde{f}_{k}$ can be recovered via integration, which is straightforward given the univariate nature of the invariant-based representation. In the following, we present a simple illustrative example that highlights the advantages of gradient-based parameterization over direct function-based approaches.
@@ 2.2.1 Monotone Univariate Networks
$h^{(\ell+1)}=\sigma\!\left(W^{(\ell)}h^{(\ell)}+b^{(\ell)}\right),\qquad\ell=0,\dots,L-1,$ (8)
with $h^{(0)}=x$, where $\sigma(\cdot)$ is the activation function and the weight matrices satisfy $W^{(\ell)}=(\widehat{W}^{(\ell)})^{2},$ for unconstrained matrices $\widehat{W}^{(\ell)}$. This construction guarantees elementwise non-negativity of the weights and, combined with the choice of activation function, ensures that $\mathcal{N}_{k}$ is monotone non-decreasing by design. A proof of the monotonicity of this neural network parameterization is provided in Appendix˜A.
@@
We consider a univariate convex function, shown in the left panel of Fig.˜1, whose derivative exhibits a stepwise structure, shown in the right panel (see Appendix˜B for details). A total of 200 samples are drawn from the corresponding gradient field, and two regression settings are considered. In the first setting, the convex function is parameterized using a convex neural network and trained such that its gradient matches the sampled data. In the second setting, the gradient is parameterized directly using a monotone neural network and trained to fit the sampled gradient observations.
@@
Using the same number of trainable parameters and identical optimizer settings, the predicted gradients for the two modeling settings are shown in Fig.˜2. The results indicate that the monotone parameterization achieves higher accuracy than the convex parameterization in fitting the sampled gradient data. This behavior is consistent across different random initializations, suggesting that the monotone formulation provides a more effective parameterization for capturing the underlying gradient structure in this example. An additional advantage of the monotone approach is improved computational efficiency during both training and inference. Because the gradient is computed directly rather than via automatic differentiation, the monotone parameterization is observed to be approximately twice as fast (see Appendix˜B).
@@
In this formulation, the basis functions are one-dimensional mappings, and thus a variety of classical function classes can be used in place of neural networks. Polynomials (in the global sense) are the simplest option, and there exist well-established methods for polynomial approximation under monotonicity constraints [52, 43, 39, 15]. However, when the underlying function exhibits multi-scale structure, as in the pedagogical example above, polynomial representations may struggle to capture sharp local variations without resorting to high polynomial degrees. Such high-degree approximations are prone to oscillatory behavior and often exhibit poor extrapolation properties. Spline-based representations provide another natural alternative, and monotonicity/convexity can be enforced through shape-restricted regression splines or constrained smoothing splines [46, 40]. In addition, monotone piecewise-cubic interpolants have been studied extensively in numerical analysis [18]. A key practical limitation of spline models, however, is that their behavior outside the data range is largely determined by boundary conditions and knot placement, and is therefore often unreliable in extrapolative regimes. In mechanics applications, constitutive laws must remain robust under loading conditions that may fall outside the calibration domain, which motivates the neural parameterizations adopted in this work.
@@ 2.3 A Probabilistic Constitutive Model via Tensorial Quantile Regression
To this end, we adopt a tensorial extension of quantile regression in which componentwise conditional stress quantiles are estimated using a structured, physics-consistent constitutive representation.
@@ 2.3.1 Tensorial Quantile Functions
$\mathrm{CDF}(\bm{P}\mid\bm{F})\triangleq\mathbb{P}(\bigwedge_{i,J}\{\bar{P}_{iJ}\leq P_{iJ}\}|\bm{F}),$ (12)
which represents the probability that all stress components lie below their specified values. This definition induces a partial order on $\mathbb{R}^{3\times 3}$ and serves as a conceptual reference for defining multivariate quantiles, rather than an object to be estimated directly.
@@
which constitutes a marginal, distribution-free predictive interval for the tensor-valued stress response. This construction does not require modeling the joint dependence structure among stress components and enables finite-sample coverage guarantees through a union-bound argument, see Appendix C.
@@ 2.3.2 Quantile Regression for Tensor-Valued Stress Prediction
$\min_{\theta}\;\mathbb{E}_{(\bm{F},\bm{P})\sim\mathcal{D}_{\mathrm{trn}}}\!\left[\sum_{i,J}\rho_{\alpha_{iJ}}\!\left(P_{iJ},q_{\alpha,iJ}(\bm{F};\theta)\right)\right],$ (17)
@@ 2.3.3 Conformal Adjustment for Distribution-Free Coverage
where the probability is taken with respect to the joint distribution of $(\bm{F},\bm{P})$. Consequently, the tensor-valued predictive set $C(\bm{F})$, defined as the Cartesian product of the componentwise conformal intervals, achieves finite-sample, distribution-free, componentwise marginal coverage; see Appendix D for the proof sketch.
@@ 3 Numerical Examples
In this section, three numerical examples are presented to demonstrate and benchmark the performance of the proposed method. In the first problem, we synthesize data from a classical constitutive model with arbitrarily chosen parameters. In the second problem, we generate synthetic data using a recently introduced constitutive model with parameters calibrated based on experimental measurements. Finally, in the third problem, we apply the proposed method directly to experimental data reported in the literature. The details of each problem are provided in the corresponding subsections, and additional information on training procedures and model parameters is given in Appendix F. For completeness, the different experimental loading conditions considered in this section are summarized in Appendix E.
@@ A Note on Experimental Loading Conditions
The conformal prediction framework described in the previous section provides distribution-free marginal coverage guarantees under the assumption of exchangeability. However, treating individual stress measurements along a loading path as exchangeable samples may lead to overly optimistic uncertainty estimates when strong intra-trajectory correlations are present. In such settings, a more appropriate exchangeability assumption is at the level of entire loading trajectories rather than individual loading steps. Accordingly, in the numerical examples presented in this work, we adopt a trajectory-level conformal calibration strategy. Specifically, for each calibration trajectory
$\mathcal{T}^{(m)}=\{(\mathbf{F}^{(m)}(t),\mathbf{P}^{(m)}(t))\}_{t=1}^{T_{m}},$
@@
$e^{(m)}(t)=\max\!\left(\hat{q}^{\mathrm{lo}}(\mathbf{F}^{(m)}(t))-\mathbf{P}^{(m)}(t),\;\mathbf{P}^{(m)}(t)-\hat{q}^{\mathrm{hi}}(\mathbf{F}^{(m)}(t)),\;0\right),$
@@
Conformal calibration is then performed across trajectories by computing the conformal quantile of the set $\{e_{\mathrm{traj}}^{(m)}\}$ over the calibration trajectories, under the assumption that the trajectories themselves are exchangeable.
@@ 3.1 Synthetic Data of the Mooney–Rivlin Model
Figure 3: Synthetic datasets generated using the Mooney–Rivlin constitutive model. (top) Twenty sets of material parameters are randomly sampled. (bottom) Mean and standard deviation computed over 500 parameter realizations.
@@ 3.1.2 Conformalized Quantile Prediction under Mixed Loading Conditions
We compare the quantile and conformalized quantile predictions for the proposed probabilistic model trained using all loading cases simultaneously. For this purpose, 15 loading trajectories are used for each loading condition as training data. The calibration set had 8 randomly selected loading trajectorirs. The other hyperparameter settings are reported in Appendix F.1. The evolution of the loss function during training is shown in Figure 5.
@@
Figure 6: (top) Quantile regression uncertainty intervals. (bottom) Conformalized quantile regression. Dashed lines denote the quantile predictions, while solid lines indicate the conformal-adjusted intervals.

### v2

# Conformal Quantile Regression for Neural Probabilistic Constitutive Modeling

@@ Abstract
Keywords Constitutive Laws $\cdot$ Physics-constrained $\cdot$ Uncertainty Quantification $\cdot$ Anisotropic Soft Materials $\cdot$ Monotonic Neural Network $\cdot$ Thermodynamics Consistency $\cdot$ Scientific Machine Learning
@@ 1 Introduction
In this work, we construct a physics-encoded, anisotropic data-driven constitutive model, building on the seminal framework of [55]. Anisotropy is incorporated through structural tensors that encode the material symmetries associated with the underlying anisotropic directions. The strain energy is represented in an additively separable invariant-based form, corresponding to a structured subclass of polyconvex energy functionals. Within this formulation, we consider neural network parameterizations for the phenomenological components of the constitutive response that preserve polyconvexity by construction. Rather than parameterizing the unknown strain energy density itself [32, 6], we parameterize its required gradients, which enter directly into the stress response. This choice leads to improved trainability and reduced computational cost.
@@
The enforcement of thermodynamic constraints in a fully Bayesian setting [12, 11] may require additional care. When physical admissibility is ensured intrinsically through architectural design (e.g., convex or polyconvex neural parameterizations), these structural properties can, in principle, be preserved under Bayesian inference. However, even in such intrinsically constrained formulations, the prior distribution must have support entirely within the admissible parameter space. For example, in convex neural networks certain weights are required to remain nonnegative to guarantee convexity; in a Bayesian formulation, this restricts the class of admissible priors to distributions with positive support (e.g., truncated Gaussian or log-normal distributions). Such restrictions can alter posterior geometry and complicate inference, particularly for variational approximations.
@@
Remark 1.1.
In this work, our primary focus is not on strictly real-time prediction, but rather on settings where data are sparse and noisy, and reliable uncertainty quantification is essential for decision-making and robust design under uncertainty. While probabilistic extensions of deterministic models often introduce additional computational cost, the proposed quantile-based approach avoids sampling during both training and inference. As a result, its computational cost at inference is comparable to that of a deterministic model.
Other Related Work
@@ 2.1 Deterministic Constitutive Backbone
We restrict attention to incompressible materials, which is a common and appropriate assumption for many soft biological and polymeric solids. Under this assumption, the third isotropic invariant is constant and therefore does not enter the strain energy representation explicitly. Specifically, the strain energy density is expressed as
$\psi(\bm{F};\{\overset{j}{\bm{G}}\}_{j=1}^{N_{G}})=\sum_{i=1}^{2}f_{i}(I_{i})+\sum_{j=1}^{N_{G}}[\bar{f}_{4j}(I_{4j})+\hat{f}_{5j}(I_{5j})],$ (1)
@@
$I_{4j}=\mathrm{tr}\![\bm{C}\,\overset{j}{\bm{G}}],\quad I_{5j}=\mathrm{tr}\![\mathrm{cof}(\bm{C})\,\overset{j}{\bm{G}}].$ (2)
@@
The functions $f_{i}$, $\bar{f}_{4j}$, and $\hat{f}_{5j}$ are unknown real-valued scalar functions. If the scalar functions are chosen to be convex with respect to their arguments, the resulting strain energy density is polyconvex in $\bm{F}$ [40, 7]. As a consequence, the associated variational problem admits minimizers under standard boundary conditions, and the resulting constitutive model satisfies rank-one convexity, which is closely related to mechanical stability. Moreover, adherence to this framework guarantees thermodynamic consistency (the Clausius–Duhem inequality is satisfied with identically zero dissipation), and satisfies the required symmetry properties with respect to the structural tensors (or fiber orientations).
@@
$\psi\!(\bm{F};\{\overset{j}{\bm{G}}\}_{j=1}^{N_{G}})=\sum_{k=1}^{M}\tilde{f}_{k}(\tilde{I}_{k}),\qquad M=2+2N_{G}.$ (3)
@@
where $\tilde{g}_{k}(\tilde{I}_{k})\equiv d\tilde{f}_{k}/d\tilde{I}_{k}$ denotes the derivative of the corresponding basis function. Since each $\tilde{f}_{k}$ is convex and non-decreasing, its derivative $\tilde{g}_{k}$ is non-negative and monotone, a property that plays a key role in the subsequent parameterization and learning strategy. For notational simplicity, the incompressibility contribution, which enters as an additional additive term $p\,\bm{F}^{-T}$, is not shown explicitly but is included in all computations.
Remark 2.1.
The representation in Equation 1 adopts an additively separable invariant-based structure, corresponding to a structured subclass of polyconvex energy functionals. Additive invariant-based forms are widely used in anisotropic constitutive modeling [25, 56, 55, 19, 8, 36, 35]. While this assumption excludes explicit cross-invariant coupling terms, it enables monotone univariate parameterizations and tractable enforcement of polyconvexity. When appropriate for the material class under consideration, the resulting structured (low-rank) representation can improve data efficiency, reduce model complexity, and facilitate post-hoc interpretability and symbolic model extraction [6].
Remark 2.2.
We note that representing anisotropy through fixed structure tensors is convenient but may be restrictive when the underlying fibre architecture is spatially varying or better described by an orientation distribution. Related distributed-orientation descriptions have been considered in earlier work by Pinsky and co-workers [50, 16].
@@ 2.2 Parameterization Strategy
Once $\tilde{g}_{k}$ has been learned, the corresponding energy functions $\tilde{f}_{k}$ can be recovered via integration, which is straightforward given the univariate nature of the invariant-based representation. In the following, we present a simple illustrative example that highlights the advantages of gradient-based parameterization over direct function-based approaches.
@@ 2.2.1 Monotone Univariate Networks
$h^{(\ell+1)}=\sigma\!(W^{(\ell)}h^{(\ell)}+b^{(\ell)}),\qquad\ell=0,\dots,L-1,$ (8)
with $h^{(0)}=x$, where $\sigma(\cdot)$ is the activation function and the weight matrices satisfy $W^{(\ell)}=(\widehat{W}^{(\ell)})^{2},$ for unconstrained matrices $\widehat{W}^{(\ell)}$. This construction guarantees elementwise non-negativity of the weights and, combined with the choice of activation function, ensures that $\mathcal{N}_{k}$ is monotone non-decreasing by design. A proof of the monotonicity of this neural network parameterization is provided in Appendix˜B.
@@
We consider a univariate convex function, shown in the left panel of Fig.˜1, whose derivative exhibits a stepwise structure, shown in the right panel (see Appendix˜C for details). A total of 200 samples are drawn from the corresponding gradient field, and two regression settings are considered. In the first setting, the convex function is parameterized using a convex neural network and trained such that its gradient matches the sampled data. In the second setting, the gradient is parameterized directly using a monotone neural network and trained to fit the sampled gradient observations.
@@
Using the same number of trainable parameters and identical optimizer settings, the predicted gradients for the two modeling settings are shown in Fig.˜2. The results indicate that the monotone parameterization achieves higher accuracy than the convex parameterization in fitting the sampled gradient data. This behavior is consistent across different random initializations, suggesting that the monotone formulation provides a more effective parameterization for capturing the underlying gradient structure in this example. An additional advantage of the monotone approach is improved computational efficiency during both training and inference. Because the gradient is computed directly rather than via automatic differentiation, the monotone parameterization is observed to be approximately twice as fast (see Appendix˜C).
@@
In this formulation, the basis functions are one-dimensional mappings, and thus a variety of function classes can be used in place of neural networks. Polynomials (in the global sense) are the simplest option, and there exist well-established methods for polynomial approximation under monotonicity constraints [58, 47, 43, 17]. However, when the underlying function exhibits multi-scale structure, as in the pedagogical example above, polynomial representations may struggle to capture sharp local variations without resorting to high polynomial degrees. Such high-degree approximations are prone to oscillatory behavior and often exhibit poor extrapolation properties. Spline-based representations provide another natural alternative, and monotonicity/convexity can be enforced through shape-restricted regression splines or constrained smoothing splines [51, 44]. In addition, monotone piecewise-cubic interpolants have been studied extensively in numerical analysis [21]. A key practical limitation of spline models, however, is that their behavior outside the data range is largely determined by boundary conditions and knot placement, and is therefore often unreliable in extrapolative regimes. In mechanics applications, constitutive laws must remain robust under loading conditions that may fall outside the calibration domain, which motivates the neural parameterizations adopted in this work.
@@ 2.3 A Probabilistic Constitutive Model via Tensorial Quantile Regression
To this end, we adopt a tensorial extension of quantile regression in which componentwise conditional stress quantiles are estimated using a structured, physics-consistent constitutive representation. A schematic representation of the overall proposed framework is shown in Fig.˜3.
Remark 2.5.
A brief summary of the probabilistic terminologies used frequently in this work is provided in Appendix A.
Figure 3: Schematic of the proposed probabilistic physics-constrained neural constitutive modeling framework. Invariant features ($I_{k}$) derived from deformation gradient $\bm{F}$ and structural tensors $\overset{j}{\bm{G}}$ are passed through monotone neural networks to learn bounds on strain-energy derivatives. These bounds induce stress predictions with quantile estimates, which are calibrated using conformal prediction for reliable uncertainty quantification.
@@ 2.3.1 Tensorial Quantile Functions
$\mathrm{CDF}(\bm{P}\mid\bm{F})\triangleq\mathbb{P}(\bigcap_{i,J}\{\bar{P}_{iJ}\leq P_{iJ}\}|\bm{F}),$ (12)
which represents the probability that all stress components lie below their specified values. Here, $\bigcap$ denotes the intersection of events. This definition induces a partial order on $\mathbb{R}^{3\times 3}$ and serves as a conceptual reference for defining multivariate quantiles, rather than an object to be estimated directly.
@@
which constitutes a marginal, distribution-free predictive interval for the tensor-valued stress response. This construction does not require modeling the joint dependence structure among stress components and enables finite-sample coverage guarantees through a union-bound argument, see Appendix D. From an engineering standpoint, the resulting quantile surfaces provide conservative, componentwise envelopes that may inform reliability-aware or worst-case design decisions. The width of these intervals also reflects the level of predictive uncertainty under the available data.
@@ 2.3.2 Quantile Regression for Tensor-Valued Stress Prediction
[…]
@@ 2.3.3 Conformal Adjustment for Distribution-Free Coverage
where the probability is taken with respect to the joint distribution of $(\bm{F},\bm{P})$. Consequently, the tensor-valued predictive set $C(\bm{F})$, defined as the Cartesian product of the componentwise conformal intervals, achieves finite-sample, distribution-free, componentwise marginal coverage; see Appendix E for the proof sketch.
Remark 2.6.
Our quantile and conformal calibration procedures are applied to each scalar stress component $P_{iJ}$ conditional on the full deformation gradient $\bm{F}$ (i.e., $\bm{F}$ is treated as the covariate and no independence between its entries is assumed). Moreover, the resulting tensor-valued predictive set is constructed as a Cartesian product of componentwise intervals; this is a conservative construction whose overall miscoverage is controlled via a union-bound argument and therefore does not require any independence assumptions among stress components (see Appendix D).
@@ 3 Numerical Examples
In this section, three numerical examples are presented to demonstrate and benchmark the performance of the proposed method. In the first problem, we synthesize data from a classical constitutive model with arbitrarily chosen parameters. In the second problem, we generate synthetic data using a recently introduced constitutive model with parameters calibrated based on experimental measurements. Finally, in the third problem, we apply the proposed method directly to experimental data reported in the literature. The details of each problem are provided in the corresponding subsections, and additional information on training procedures and model parameters is given in Appendix G. For completeness, the different experimental loading conditions considered in this section are summarized in Appendix F.
@@ A Note on Experimental Loading Conditions
The conformal prediction framework described in the previous section provides distribution-free marginal coverage guarantees under the assumption of exchangeability. A natural interpretation of this assumption is a pooled (pointwise) calibration strategy, in which all observations across trajectories are treated as exchangeable samples and nonconformity scores are computed at the level of individual loading steps. However, this assumption may be inappropriate in the presence of strong intra-trajectory correlations. When data are collected along structured, prescribed loading paths (e.g., a monotonically increasing uniaxial test), treating individual stress measurements as exchangeable samples can lead to overly optimistic (i.e., under-covered) uncertainty estimates. In such settings, a more appropriate exchangeability assumption is at the level of entire loading trajectories.
Accordingly, in the numerical examples presented in this work, we adopt a trajectory-level conformal calibration strategy due to the use of data generated from structured loading paths. Specifically, for each calibration trajectory
$\mathcal{T}^{(m)}=\{(\bm{F}^{(m)}(t),\bm{P}^{(m)}(t))\}_{t=1}^{T_{m}},$
@@
$e^{(m)}(t)=\max\!\left(\hat{q}^{\mathrm{lo}}(\bm{F}^{(m)}(t))-\bm{P}^{(m)}(t),\;\bm{P}^{(m)}(t)-\hat{q}^{\mathrm{hi}}(\bm{F}^{(m)}(t)),\;0\right),$
@@
This trajectory-level formulation applies the conformal adjustment defined in Section 2.3.3 to the aggregated scores $\{e_{\mathrm{traj}}^{(m)}\}$, rather than to individual pointwise scores. Conformal calibration is then performed across trajectories by computing the conformal quantile of the set $\{e_{\mathrm{traj}}^{(m)}\}$ over the calibration trajectories, under the assumption that the trajectories themselves are exchangeable.
@@ 3.1 Synthetic Data of the Mooney–Rivlin Model
Figure 4: Synthetic datasets generated using the Mooney–Rivlin constitutive model. (top) Twenty sets of material parameters are randomly sampled, with black dots indicating the corresponding data points. (bottom) Mean and standard deviation computed over 500 parameter realizations.
@@ 3.1.2 Conformalized Quantile Prediction under Mixed Loading Conditions
We compare the quantile and conformalized quantile predictions for the proposed probabilistic model trained using all loading cases simultaneously. For this purpose, 15 loading trajectories are used for each loading condition as training data. The calibration set had 8 randomly selected loading trajectorirs. The other hyperparameter settings are reported in Appendix G.1. The evolution of the loss function during training is shown in Figure 6.
@@
Figure 7: (top) Quantile regression uncertainty intervals. (bottom) Conformalized quantile regression. Dashed lines denote the quantile predictions, while solid lines indicate the conformal-adjusted intervals. Data points are shown as black dots.